if the sum of a number and 7 is tripled the result is 3 less than twice the number
Answer by tutorcecilia(2144)

if the sum of a number and 7 is tripled the result is 3 less than twice the number
Let x = "a number"
Sum = to add
Result = "equal"
Tripled = times 3
Less than 3 = substract 3
Twice = times 2
.
3(x+7)=2x-3
3x+21=2x-3 [isolate the variable (x)]
3x-2x+21=2x-2x-3
x+21=-3
x+21-21=-3-21
x=-24
.
Check by plugging (x=-24) back into the original equation:
3(x+7)=2x-3
3(-24+7)=2(-24)-3
3(-17)=-48-3
-51=-51